Qd = 180 - 2P,

Qs = ‐ 15 + P

The market is government-regulated with a price support per unit and production quotas. If the price is set at $72 per unit, what production quota is needed to make sure there are no shortages or surpluses?

 

a.an increase in the demand for the good. new demand equation is Qd = 190 - 2P. The government is trying to decide between two options:

Maintain the number of quotas and let the market adjust, or

Maintain the price support and increase the number of quotas.

 



1
Expert's answer
2021-10-18T11:29:57-0400
"180-2p=-15+p"

"3p=195"

"p=65"

If p=72, then


"Q_D=180-2\times72=36"

"Q_S=-15+72=57"

"\varDelta Q=57-36=21"

If


"Q_D=190-2p"

"190-2p=-15+p"

"3p=205"

"p=68.3"

"Q_D=180-2\times68.3=43.4"

"Q_S=-15+68.3=53.3"

"\varDelta Q=53.3-43.4=9.9"